# Do Language Models Have a Critical Period for Language Acquisition?

### v1

@@ Abstract
Humans appear to have a critical period (CP) for language acquisition: Second language ($\mathrm{L}_{2}$) acquisition becomes harder after early childhood, and ceasing exposure to a first language ($\mathrm{L}_{1}$) after this period (but not before) typically does not lead to substantial loss of $\mathrm{L}_{1}$ proficiency. It is unknown whether these CP effects result from innately determined brain maturation or as a stabilization of neural connections naturally induced by experience. In this study, we use language models (LMs) to test the extent to which these phenomena are peculiar to humans, or shared by a broader class of language learners. We vary the age of exposure by training LMs on language pairs in various experimental conditions, and find that LMs, which lack any direct analog to innate maturational stages, do not show CP effects when trained sequentially on $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$. Our results contradict the claim that CP effects are an inevitable result of learning in statistical learners, and they are consistent with an innate mechanism for CP effects. We show that we can reverse-engineer the CP by introducing a regularizer partway through training to simulate a maturational decrease in plasticity. All in all, our results suggest that $\mathrm{L}_{1}$ learning on its own may not be enough to induce a CP, and additional engineering is necessary to make language models more cognitively plausible.
@@ 1 Introduction
The tension between nature and nurture is central to the question of how humans acquire language. The Critical Period (CP) for language acquisition is no exception. Around the onset of adolescence, humans exhibit a loss in ability to acquire a second language through immersion and a tendency not to forget their first language under deprivation (Penfield and Roberts, 1959; Lenneberg, 1967; Johnson and Newport, 1989; Pallier et al., 2003). Scholars of human development have long debated whether these phenomena are predetermined by innately encoded developmental changes in the maturing brain (Penfield, 1965; Chomsky, 1965; Pinker, 1994), or natural consequences of increased experience that any typical statistical learner would be subject to (Elman et al., 1996; Seidenberg and Zevin, 2006; Thiessen et al., 2016).
@@
Until recently, it was difficult to differentiate these two hypotheses; as we could only observe one kind of statistical learner (i.e., humans), we could not identify which properties of its learning process were responsible for its behavior (Warstadt and Bowman, 2022). Recent improvement in neural language modeling upends this state of affairs: Language Models (LMs) can learn to simulate many native-like grammatical judgments (Warstadt et al., 2020; Zhang et al., 2021; Hu et al., 2020)—long regarded as one of the main behavioral measures of native speaker knowledge (Chomsky, 1957; Johnson and Newport, 1989)—and, like humans, they acquire this knowledge from unstructured input without the need for negative evidence. However, their learning patterns differ from those of humans in a number of ways. LMs can thus provide additional information about which phenomena are likely to be typical of general language learners, and which are peculiar to humans.
@@
In this work, we use language models to study the CP for language acquisition, focusing on second language acquisition and first language attrition. Our experiments test for CP effects [^1] While the term “critical period“ is sometimes used to refer to a specific biological construct, we use it pretheoretically to refer to the characteristic effects of age-of-exposure on $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ performance. For clarity, we prefer to use the term “critical period effects”, but we intend these two terms interchangeably. in LMs by training them from scratinnate ch on bilingual data, varying only the age of exposure to $\mathrm{L}_{2}$. We test whether LMs learn $\mathrm{L}_{2}$ more easily when exposed to it simultaneously with $\mathrm{L}_{1}$, or when exposed to it only after learning $\mathrm{L}_{1}$; similarly, we test whether they forget or not $\mathrm{L}_{1}$ after extensive training on it. Experimentally, we find that LMs are unlike humans in both respects, showing neither of the CP effects. Thus, our results contradict the view that CP effects are an expected consequence of statistical learner, and they are consistent with (but only provide weak evidence for) the view that the CP in humans is a biologically programmed developmental stage.
@@
The benefits of studying the CP in neural models, however, go beyond just discerning the two hypotheses above. If LMs do differ from humans, it may be useful to attempt to reverse-engineer those learning properties exhibited by humans (Dupoux, 2018). Furthermore, minimizing differences between LMs and humans is a necessary step in enabling their use as models of human language acquisition (Warstadt and Bowman, 2022). Thus, we also attempt to reverse-engineer a CP by simulating a loss of neural plasticity using Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), a Bayesian regularizer used in machine learning to mitigate catastrophic forgetting. Our experiments show that both of the CP effects emerge in tandem when the model’s plasticity is explicitly controlled in this way. Our findings demonstrate the utility of LMs as tools for theories about human language acquisition, and they suggest a path forward to making LMs more developmentally plausible models of human language acquisition.
@@ 2 Background: The Critical Period
The idea that there is a critical period (or sensitive period) for language learning has long been prominent in language acquisition research (Penfield and Roberts, 1959; Lenneberg, 1967). Discussions around the CP, however, typically cluster a number of related observations; these must be teased apart in order to be properly understood (Singleton, 2005; Mayberry and Kluender, 2018). The critical period can be divided into 3 main phenomena, these regard: a CP for $\mathrm{L}_{1}$ acquisition, a CP for $\mathrm{L}_{2}$ acquisition, and a CP for $\mathrm{L}_{1}$ attrition. We focus on the latter two here.222 There is also clear evidence of a CP for $\mathrm{L}_{1}$ acquisition, as late $\mathrm{L}_{1}$ acquisition in the deaf community is not uncommon (Mayberry and Fischer, 1989; Newport, 1990). However, simulating late $\mathrm{L}_{1}$ exposure requires training on non-linguistic data, which is beyond the scope of this work.
@@ 2.1 Critical Period for $\mathrm{L}_{2}$ Acquisition
CP effects for $\mathrm{L}_{2}$ acquisition consist of greater difficulty in learning a second language and worse learning outcomes as the age of exposure increases. As the beginning of $\mathrm{L}_{2}$ exposure is highly variable, this is perhaps the most well-known of the CP phenomena. The effects of age of exposure on phonological performance (i.e., one’s accent) is part of folk knowledge, and was a key piece of evidence in the first works to propose a neurological mechanism for these effects (Lenneberg, 1967). Numerous studies also show that age of exposure correlates with worse $\mathrm{L}_{2}$ performance on morphological and syntactic acceptability judgment tasks (Johnson and Newport, 1989; Hartshorne et al., 2018). While the exact nature and reliability of these effects has been questioned at times (Ioup et al., 1994), the existence of age-of-exposure effects is generally accepted. We refer the reader to several thorough reviews of the relevant evidence (Mayberry and Kluender, 2018; Thiessen et al., 2016; Singleton, 2005).
In the realm of non-human learning, no prior work has tested this CP in a controlled way. In a more general form, however, $\mathrm{L}_{2}$ acquisition has been studied in depth (Dufter and Schütze, 2020; Chen et al., 2023, inter alia). Most related to our work, Oba et al. (2023) trained a number of language models on an $\mathrm{L}_{1}$ and then fine-tuned these models on both $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$; they find that, unlike humans, this two-step training improves LMs’ $\mathrm{L}_{2}$ performance. However, they fine-tune their models on $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ simultaneously within the same bidirectional transformer context, which we contend is not cognitively plausible.
Beyond Oba et al.’s (2023) study, weak evidence against the existence of a CP for $\mathrm{L}_{2}$ on neural networks is suggested by a large body of work which fine-tunes pre-trained neural networks to perform new tasks (e.g., Devlin et al., 2019; Liu et al., 2019; Driess et al., 2023). These studies indicate that a neural model can achieve comparable, or sometimes even superior, performance on a fine-tuned task compared to training on it from scratch. However, their main focus in on tasks other than language modeling itself, such as classification.
@@ 2.2 Critical Period for $\mathrm{L}_{1}$ Attrition
The CP for $\mathrm{L}_{1}$ attrition refers to a loss of proficiency in $\mathrm{L}_{1}$ due to a lack of exposure to it. This phenomenon is largely constrained to earlier ages (Pallier, 2007), as adults who emigrate from their $\mathrm{L}_{1}$ community do not typically forget $\mathrm{L}_{1}$ entirely. However, profound attrition is possible if $\mathrm{L}_{1}$ exposure ceases during childhood. For example, Pallier et al. (2003) studied Korean-born adoptees in France who had no recognition of their $\mathrm{L}_{1}$, despite living in Korea for as long as eight years.
In the computational domain, language attrition relates to another large body of work in life-long and continual learning. In short, a large number of works have shown that neural networks are prone to catastrophic forgetting (McCloskey and Cohen, 1989; French, 1999), losing most of their proficiency in their original training domain when fine-tuned on another. To mitigate catastrophic forgetting, continual learning aims to improve neural networks by providing them the ability to retain earlier proficiencies. This is done in multiple ways: through the use of adapters, regularizers (Kirkpatrick et al., 2017; Pan et al., 2020), or further training in the original domain.
@@ 2.3 Theories
Innate accounts of the CP argue that this loss in plasticity is driven by properties which are specific to how humans acquire language. Some of these accounts are based on the hypothesis that children—but not adults—are equipped with a specialized language acquisition device such as Universal Grammar (Chomsky, 1965; Newport, 1990). On this view, the CP occurs when Universal Grammar is (wholly or partially) lost, displaced, or dismantled as we age, which would explain why adults struggle with language acquisition (Chomsky, 1965, p. 207; Borer and Wexler, 1987; Bley-Vroman et al., 1988; Schachter, 1988; Pinker, 1994, p. 294). [^4] Child and adult language acquisition are seen as driven by qualitatively different mechanisms on this view (Thiessen et al., 2016), which is supported by evidence that general analytic ability predicts adult—but not child—$\mathrm{L}_{2}$ learning outcomes (DeKeyser, 2000). Other innate accounts are not language-specific, especially those with an explicit neurobiological basis. For example, humans (and other mammals; Paolicelli et al., 2011) go through a phase of synaptic pruning peaking in late childhood and adolescence (Huttenlocher, 1979, 1990) during which disused neuronal connections are reduced (Hensch, 2005). Monolingual brains show signs of more extensive pruning than bilingual ones (Mechelli et al., 2004), suggesting that early in life abundant synapses provide the necessary plasticity to acquire a second language with ease, and that later these synapses may be pruned if not yet recruited (De Bot, 2006). Beyond synaptic pruning, other neurobiological process such as myelination of (Pulvermüller and Schumann, 1994; Pujol et al., 2006) and lateralization (Lenneberg, 1967) are also correlated with a loss in plasticity as we age. [^5] Often, these processes have experiential correlates as well, i.e., their outcomes are modulated by experiences during development (Mechelli et al., 2004; Cheng et al., 2019). This is consistent, however, with the timing and onset of the process being biologically determined.
@@
By contrast, experiential accounts of the CP argue that a loss in plasticity is a consequence of learning itself (Munro, 1986; Elman et al., 1996, p. 283; Ellis and Lambon Ralph, 2000; Zevin and Seidenberg, 2002; Seidenberg and Zevin, 2006; Thiessen et al., 2016; Achille et al., 2019). Early experiments on connectionist models found that stages associated with human development sometimes fall out naturally during the training of low-bias neural networks (McClelland, 1989). Connectionist word learning simulations found an effect of age-of-acquisition on learning outcomes (Ellis and Lambon Ralph, 2000; Zevin and Seidenberg, 2002). Numerous scholars explain this loss of plasticity—sometimes referred to as entrenchment—as a natural consequence of the training dynamics of neural networks that lead to convergence (Munro, 1986; Elman et al., 1996; Ellis and Lambon Ralph, 2000; Seidenberg and Zevin, 2006). As Ellis and Lambon Ralph (2000, p. 1108) argue, in a model with random weights (e.g., shortly after initialization) the activations of individual units tend towards intermediate values, leading to large weight changes, but as training proceeds, the units’ activations tend towards extreme values making them less prone to change, even if the prediction loss is large.
@@ 3 The Role of LMs in Studying the CP
Computational models have the potential to be an irreplaceable tool for informing debates about language acquisition, as they enable a degree of control over the learning mechanism and environment not possible with human subjects; their relevance to questions about human learning, however, is hampered by their numerous differences from human learners (Warstadt and Bowman, 2022). Nonetheless, there are some theoretical claims that current LMs can provide strong or even conclusive evidence about. Not surprisingly, these models are increasingly being used to test theories of language acquisition (McCoy et al., 2020; Lavechin et al., 2021; Wilcox et al., 2023; Warstadt et al., 2023). Language models can, for example, refute some poverty of the stimulus claims by providing existence proofs about language learnability (Clark and Lappin, 2011, p. 30).
@@
The strong innate claim is implicit in the argument that the mere existence of CP effects counts as evidence in favor of an innate mechanism like Universal Grammar (see, e.g.  Schachter, 1988). This argument depends on the premise that a change in learning ability as extreme as what is seen in $\mathrm{L}_{2}$ acquisition could not (or would be very unlikely to) arise from a single domain-general learning mechanism. This premise, and thus the argument, is simple to refute by finding a counterexample, that is an instance of a low-bias learner that does show CP effects. Transformer-based LMs, while not bias-free, have proven to be effective learners for vision (Dosovitskiy et al., 2021), protein folding (Jumper et al., 2021), and many other types of data, suggesting that they are sufficiently domain-general to refute strong innate claim if they do innately show CP effects.

### v2

# Investigating Critical Period Effects in Language Acquisition through Neural Language Models

@@ Abstract
Humans appear to have a critical period (CP) for language acquisition: Second language ($\mathrm{L}_{2}$) acquisition becomes harder after early childhood, and ceasing exposure to a first language ($\mathrm{L}_{1}$) after this period (but not before) typically does not lead to substantial loss of $\mathrm{L}_{1}$ proficiency. It is unknown whether these CP effects result from innately determined brain maturation or as a stabilization of neural connections naturally induced by experience. In this study, we use language models (LMs) to test the extent to which these phenomena are peculiar to humans, or shared by a broader class of language learners. We vary the age of exposure by training LMs on language pairs in various experimental conditions, and find that LMs, which lack any direct analog to innate maturational stages, do not show CP effects when the age of exposure of $\mathrm{L}_{2}$ is delayed. Our results contradict the claim that CP effects are an inevitable result of statistical learning, and they are consistent with an innate mechanism for CP effects. We show that we can reverse-engineer the CP by introducing a regularizer partway through training to simulate a maturational decrease in plasticity. All in all, our results suggest that $\mathrm{L}_{1}$ learning on its own may not be enough to induce a CP, and additional engineering is necessary to make language models more cognitively plausible.
@@ 1 Introduction
The tension between nature and nurture is central to questions surrounding how humans acquire language. The Critical Period (CP) for language acquisition is no exception. Around the onset of adolescence, humans exhibit a loss in ability to acquire a second language through immersion and a tendency not to forget their first language under deprivation (Penfield and Roberts, 1959; Lenneberg, 1967; Johnson and Newport, 1989; Pallier et al., 2003). Scholars of human development have long debated whether these phenomena are predetermined by innately encoded developmental changes in the maturing brain (Penfield, 1965; Chomsky, 1965; Pinker, 1994), or natural consequences of increased experience that any typical statistical learner would be subject to (Elman et al., 1996; Seidenberg and Zevin, 2006; Thiessen et al., 2016).
@@
Until recently, it was difficult to differentiate these two hypotheses; as we could only observe one kind of statistical learner (i.e., humans), we could not identify which properties of its learning process were responsible for its behavior. Recent improvement in neural language modeling upends this state of affairs (Warstadt and Bowman, 2022). Language Models (LMs) can learn to simulate many native-like grammatical judgments (Warstadt et al., 2020; Zhang et al., 2021; Hu et al., 2020)—long regarded as one of the main behavioral measures of native speaker knowledge (Chomsky, 1957; Johnson and Newport, 1989)—and, like humans, they acquire this knowledge from unstructured input without the need for negative evidence. However, their learning algorithm and structure differ from those of humans in a number of ways. LMs can thus provide additional information about which phenomena are likely to be typical of general language learners, and which are peculiar to humans.
@@
In this work, we use language models to study the CP for language acquisition, focusing on second language acquisition and first language attrition. Our experiments test for CP effects [^1] The term critical period is sometimes used to refer to a specific biological construct. For clarity, we use the term CP effects to refer to the characteristic observable effects of age of exposure on $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ performance. in LMs by training them from scratch on bilingual data, varying only the age of exposure to $\mathrm{L}_{2}$. We test whether, like humans, LMs learn $\mathrm{L}_{2}$ more easily when exposed to it simultaneously with $\mathrm{L}_{1}$ from the beginning of training, rather than when exposed to it only after learning $\mathrm{L}_{1}$. Similarly, we test whether they fail to forget $\mathrm{L}_{1}$ after extensive training on it. Experimentally, we find that LMs are unlike humans in both respects.222 We note that our results thus align with the well-known fact that language models: (i) are prone to catastrophic forgetting; and (ii) are good at transfer learning. Our experiments, though, support the novel conclusion that transfer performance (training in $\mathrm{L}_{1}$ followed by $\mathrm{L}_{2}$) leads to similar or better results than jointly training on both languages. Thus, our results contradict the view that CP effects are an expected consequence of statistical learning, and they are consistent with (but only provide weak evidence for) the view that the CP in humans is a biologically programmed developmental stage.
@@
The benefits of studying the CP in neural models, however, go beyond just discerning the two hypotheses above. If LMs do differ from humans, it may be useful to attempt to reverse-engineer those learning properties exhibited by humans (Dupoux, 2016). Furthermore, minimizing differences between LMs and humans is a necessary step in enabling their use more broadly as models of human language acquisition (Warstadt and Bowman, 2022). Thus, we also attempt to reverse-engineer a CP by simulating a loss of neural plasticity using Elastic Weight Consolidation (EWC; Kirkpatrick et al., 2017), a Bayesian regularizer used in machine learning to mitigate catastrophic forgetting. Our experiments show that both of the CP effects emerge in tandem when the model’s plasticity is explicitly controlled in this way. Our findings demonstrate the utility of LMs as tools for theories about human language acquisition, and they suggest a path forward to making LMs more developmentally plausible models of human language acquisition.
@@ 2 Background: The Critical Period
The proposition that there is a critical period for language learning has long been prominent in language acquisition research (Penfield and Roberts, 1959; Lenneberg, 1967). Discussions around the CP, however, typically cluster a number of related observations; these must be teased apart in order to be properly understood (Singleton, 2005; Mayberry and Kluender, 2018). The critical period can be divided into 3 main phenomena, namely: a CP for $\mathrm{L}_{1}$ acquisition, a CP for $\mathrm{L}_{2}$ acquisition, and a CP for $\mathrm{L}_{1}$ attrition. We focus on the latter two here.333 Strong evidence of a CP for $\mathrm{L}_{1}$ acquisition has been demonstrated in late $\mathrm{L}_{1}$ learners from the deaf community (Mayberry and Fischer, 1989; Newport, 1990). However, as simulating late $\mathrm{L}_{1}$ exposure requires training on non-linguistic data, we consider it beyond the scope of this work.
@@ 2.1 Critical Period for $\mathrm{L}_{2}$ Acquisition
CP effects for $\mathrm{L}_{2}$ acquisition consist of greater difficulty in learning a second language and worse learning outcomes as the age of exposure increases. As humans vary greatly in the beginning of $\mathrm{L}_{2}$ exposure, this is perhaps the most well-known of the CP phenomena. The effects of age of exposure on phonetics and phonology (i.e., one’s accent) are part of folk knowledge and were a key piece of evidence in the first works to propose a neurological mechanism for these (Lenneberg, 1967). Numerous studies also show that age of exposure correlates with worse $\mathrm{L}_{2}$ performance on morphological and syntactic acceptability judgment tasks (Johnson and Newport, 1989; Hartshorne et al., 2018). While the exact nature and reliability of these effects has been questioned at times (Ioup et al., 1994), the existence of age-of-exposure effects is generally accepted. We refer the reader to several thorough reviews of the relevant evidence (Singleton, 2005; Thiessen et al., 2016; Mayberry and Kluender, 2018).
In the realm of computational learners, no prior work has tested this CP in a controlled manner. In a more general form, however, $\mathrm{L}_{2}$ acquisition has been studied in depth (Dufter and Schütze, 2020; Chen et al., 2023, inter alia). Most related to our work, Oba et al. (2023) trained a number of language models on an $\mathrm{L}_{1}$ and then fine-tuned these models on both $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$; they find that, unlike humans, this two-step training improves LMs’ $\mathrm{L}_{2}$ performance. This already suggests that LMs may not show CP effects for $\mathrm{L}_{2}$ learning. However, they do make $\mathrm{L}_{2}$ learning relatively easier by fine-tuning their models on $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ simultaneously within the same bidirectional transformer context, which we contend is not cognitively plausible.
Beyond Oba et al.’s (2023) study, weak evidence against the existence of a CP for $\mathrm{L}_{2}$ in neural networks is suggested by a large body of work on transfer learning which fine-tunes pretrained neural networks to perform new tasks (e.g., Devlin et al., 2019; Liu et al., 2019; Driess et al., 2023). These studies indicate that a neural model can achieve superior performance on a fine-tuned task compared to training on it from scratch. However, these studies do not manipulate age-of-exposure while controlling for the total amount of input, and their main focus is on tasks other than language modeling itself, such as classification.
@@ 2.2 Critical Period for $\mathrm{L}_{1}$ Attrition
The CP for $\mathrm{L}_{1}$ attrition refers to a loss of proficiency in $\mathrm{L}_{1}$ due to a lack of exposure to it. This phenomenon is largely constrained to earlier ages (Pallier, 2007), as adults who emigrate from their $\mathrm{L}_{1}$ community do not typically forget their $\mathrm{L}_{1}$ entirely. However, profound attrition is possible if $\mathrm{L}_{1}$ exposure ceases during childhood. For example, Pallier et al. (2003) studied Korean-born adoptees in France who had no recognition of their $\mathrm{L}_{1}$, despite living in Korea for as long as eight years.
In the computational domain, language attrition relates to another large body of work in life-long and continual learning. In short, a large number of works have shown that neural networks are prone to catastrophic forgetting (McCloskey and Cohen, 1989; French, 1999), losing most of their proficiency in their original training domain when fine-tuned on another. Continual learning mitigates catastrophic forgetting through the use of adapters (Houlsby et al., 2019; Pfeiffer et al., 2020), regularizers (Kirkpatrick et al., 2017; Pan et al., 2020), or further training in the original domain.
@@ 2.3 Theories
Innate accounts of the CP argue that this loss in plasticity is driven by properties which are specific to how humans acquire language. Some of these accounts are based on the hypothesis that children—but not adults—are equipped with a specialized language acquisition device such as Universal Grammar (Chomsky, 1965; Newport, 1990). On this view, the CP occurs when Universal Grammar is (wholly or partially) lost, displaced, or dismantled as we age, which would explain why adults struggle with language acquisition (Chomsky, 1965, p. 207; Borer and Wexler, 1987; Bley-Vroman et al., 1988; Schachter, 1988; Pinker, 1994, p. 294). [^5] Child and adult language acquisition are seen as driven by different mechanisms on this view (Thiessen et al., 2016), supported by evidence that general analytic ability predicts adult—but not child—$\mathrm{L}_{2}$ learning outcomes (DeKeyser, 2000). Other innate accounts are not language-specific, especially those with an explicit neurobiological basis. For example, humans (and other mammals; Paolicelli et al., 2011) go through a phase of synaptic pruning peaking in late childhood and adolescence (Huttenlocher, 1979, 1990) during which disused neuronal connections are reduced (Hensch, 2005). Monolingual brains show signs of more extensive pruning than bilingual ones (Mechelli et al., 2004), suggesting that early in life abundant synapses provide the necessary plasticity to acquire a second language with ease, and that later these synapses may be pruned if not yet recruited (De Bot, 2006). Beyond synaptic pruning, other neurobiological process such as myelination (Pulvermüller and Schumann, 1994; Pujol et al., 2006) and lateralization (Lenneberg, 1967) are also correlated with a loss in plasticity as we age. [^6] Often, these processes have experiential correlates as well, i.e., their outcomes are modulated by experiences during development (Mechelli et al., 2004; Cheng et al., 2019). This is consistent, however, with the timing and onset of the process being biologically determined.
@@
By contrast, experiential accounts of the CP argue that a loss in plasticity is a consequence of learning itself (Munro, 1986; Elman et al., 1996, p. 283; Ellis and Lambon Ralph, 2000; Zevin and Seidenberg, 2002; Seidenberg and Zevin, 2006; Thiessen et al., 2016; Achille et al., 2019). Early experiments on connectionist models found that stages associated with human development sometimes fall out naturally during the training of low-bias neural networks (McClelland, 1989). Connectionist word learning simulations found an effect of age of acquisition on learning outcomes (Ellis and Lambon Ralph, 2000; Zevin and Seidenberg, 2002). Numerous scholars explain this loss of plasticity—sometimes referred to as entrenchment—as a natural consequence of the training dynamics of networks that lead to convergence (Munro, 1986; Elman et al., 1996; Ellis and Lambon Ralph, 2000; Seidenberg and Zevin, 2006). As Ellis and Lambon Ralph (2000, p. 1108) argue, in a model with random weights (e.g., after initialization) the activations of individual units tend towards intermediate values, leading to large weight changes, but as training proceeds, the units’ activations tend towards extreme values making them less prone to change, even if the prediction loss is large.
@@ 3 The Role of LMs in Studying the CP
Computational models have the potential to be a powerful tool for informing debates about language acquisition, as they enable a degree of control over the learning mechanism and environment not possible with human subjects; their relevance to questions about human learning, however, is hampered by their numerous differences from human learners (Warstadt and Bowman, 2022). Nonetheless, there are some theoretical claims that current LMs can provide strong or even conclusive evidence about. Not surprisingly, these models are increasingly being used to test theories of language acquisition (McCoy et al., 2020; Lavechin et al., 2021; Wilcox et al., 2023; Warstadt et al., 2023). Language models can, for example, refute some poverty of the stimulus claims by providing existence proofs about language learnability (Clark and Lappin, 2011, p. 30).
@@
The strong innate claim is implicit in the argument that the mere existence of CP effects counts as evidence in favor of an innate mechanism like Universal Grammar (see, e.g.,  Schachter, 1988). This argument depends on the premise that a change in learning ability as extreme as what is seen in $\mathrm{L}_{2}$ acquisition could not (or would be very unlikely to) arise from a single domain-general learning mechanism. This premise, and thus the argument, is simple to refute by finding a counterexample, that is, an instance of a low-bias learner that does show CP effects. Transformer-based LMs, while not bias-free, have proven to be effective learners for vision (Dosovitskiy et al., 2021), protein folding (Jumper et al., 2021), and many other types of data, suggesting that they are sufficiently domain-general to refute strong innate claim if they do show CP effects.